OpenStudy (bekkah323):

the point (_,0) is the center of the circle (x-2)^2+y^2=16

OpenStudy (anonymous):

i am willing to bet you can guess the answer

OpenStudy (bekkah323):

?

OpenStudy (anonymous):

make a guess. pick one of the numbers you are looking at. don't pick 16

OpenStudy (bekkah323):

2?

OpenStudy (anonymous):

you win

OpenStudy (bekkah323):

but why?

OpenStudy (anonymous):

16 is the square of the radius, so the radius if 4

OpenStudy (anonymous):

why is a good question

OpenStudy (anonymous):

general form of a circle is \[(x-\color{red}h)^2+(y-\color{blue}k)^2=r^2\] for a circle with center \((\color{red}h,\color{blue}k)\) and radius \(r\) it comes from the distance formula

OpenStudy (anonymous):

you have \[(x-2)^2+(y-0)^2=4^2\] so your radius is 4, and the center is \((2,0)\)

OpenStudy (bekkah323):

ok thanks, that makes sense

OpenStudy (anonymous):

yw